# Orientable and negative orientable sequences

### v1

@@ 3.3 Construction II: A construction for general $n$
We first need the following definition.
@@
It remains to establish how to construct the desired sequences $S$. Suppose, for $q>2$ and $n\geq 2$, $G$ is a directed graph with vertices the $q$-ary ${(n-1)}$-tuples with pseudoweight less than $nq/2-1$, and with a directed edge connecting vertices $\mathbf{u}=(u_{0},u_{1},\ldots,u_{n-2})$ and $\mathbf{v}=(v_{0},v_{1},\ldots,v_{n-2})$ if and only if $u_{i}=v_{i-1}$, $(1\leq i\leq n-2)$, and $w^{*}(\mathbf{u})+f(v_{n-2})<nq/2$ (where $f$ is as in Definition 3.13). Clearly $G$ is a subgraph of the de Bruijn digraph. As is conventional, we identify the edge connecting $(u_{0},u_{1},\ldots,u_{n-2})$ to $(u_{1},u_{2},\ldots,u_{n-2},v)$ with the $n$-tuple $(u_{0},u_{1},\ldots,u_{n-2},v)$, and $G$ contains as an edge every $q$-ary $n$-tuple with pseudoweight at most $nq/2$.
@@
must satisfy $t+w^{*}(\mathbf{u})<nq/2$. That is, the in-degree of every vertex is the same as its out-degree. Moreover, $G$ is clearly connected, as it is straightforward to construct a ‘low negaweight’ directed path between any pair of vertices — that is, for any vertex $\mathbf{u}=(u_{0},u_{1},\ldots,u_{n-2})$ there is a path to $(1,1,\ldots,1)$ via vertices of the form $(u_{i},u_{i+1},\ldots,u_{n-2},1,1,\ldots,1)$ and similarly from $(1,1,\ldots,1)$ to any vertex $\mathbf{v}$. This means there exists a directed Eulerian Circuit in $G$ (see, for example, [10, Corollary 6.1]), yielding a $q$-ary $n$-window sequence $S$ with the property that all the $n$-tuples appearing in $S$ have weight less than $nq/2$, and with period equal to the number of $q$-ary $n$-tuples with pseudoweight less than $nq/2$.
The above discussion establishes the following result.
@@
There exists a $\mathcal{NOS}_{q}(n)$ of period $\frac{q^{n}-r_{q,n,nq/2}}{2}$ for all $q>2$ and $n\geq 2$, where $r_{q,n,s}$ denotes the number of $q$-ary $n$-tuples with pseudoweight exactly $s$.
@@ 3.4 Enumerating $n$-tuples with given pseudoweight
Lemma 3.15 means that it is of interest to know the value of $r_{q,n,s}$. Observe that if $q$ is even $r_{q,n,s}$ is only defined for integer values $s$, and if $q$ is odd it is only defined when $s$ is a multiple of 0.5.
We have the following elementary result.
@@
As above, suppose $r_{q,n,s}$ ($q\geq 2$, $n\geq 1$) is the number of $q$-ary $n$-tuples with pseudoweight exactly $s$, where $r_{q,n,s}=0$ by definition if $s<n$ or $s>n(q-1)$.
@@
    Parts (i)–(iv) are trivially true (observing that the 1-tuples $(0)$ and $(q/2)$ both have pseudoweight $q/2$). Part (v) follows immediately by considering the $q$ possible ways of adding a final entry to an $(n-1)$-tuple to make an $n$-tuple of a given pseudoweight, observing that inserting 00 or $q/2$ as the final entry will increase the pseudoweight by $q/2$.
The recursions of Lemma 3.16 make calculating the set of values $r_{q,n,s}$ straightforward for small $n$ and $q$. We can also make some simple observations.
@@
    if $n\geq 1$ and $0\leq i\leq 2n$ then $r_{3,n,n+i/2}=N_{2}(i,n)$, where $N_{2}(i,n)$ is a trinomial coefficient (see Section 3.7 below);
@@
      If $w\neq q$, then, by the same result $r_{q,2,w}$ will be the sum of $q$ terms at most $q-1$ of which are equal to 1, i.e. $r_{q,2,w}\leq q-1$.
@@
      If $w\neq q$, then, by the same result $r_{q,2,w}$ is equal to the sum of $q-1$ values (at most $q-2$ of which are non-zero) and $r_{q,n-1,w-q/2}$ (which is at most one since $w\neq q/2$). All but at most one of the $q-2$ non-zero values are equal to one (the remaining value being equal to two). Hence $r_{q,2,w}\leq(q-3)+2+1=q$.
@@
      This follows by induction on $n$. If $n=3$ then, by Lemma 3.16(v), $r_{q,3,w}$ is the sum of $q$ values, all of which are less than $q$ (apart from at most one value equal to $q$) — from (i). Hence, since $q>2$, $r_{q,3,w}\leq(q-1)^{2}+q<q^{2}$, and the claim holds. The induction step is trivial since $r_{q,m+1,w}$ is equal to the sum of $q$ values $r_{q,m,i}$, each of which is less than $q^{m-1}$.
@@
      This also follows by induction on $n$. If $n=3$ then, by Lemma 3.16(v), $r_{q,3,w}$ is the sum of $q$ values, all of which are at most $q$ (apart from at most one value equal to $q+2$) — from (ii). Hence, $r_{q,3,w}\leq(q-1)q+(q+2)=q^{2}+2$, and the claim holds. The induction step is trivial since $r_{q,m+1,w}$ is equal to the sum of $q$ values $r_{q,m,i}$, each of which is less than $q^{m-3}(q^{2}+2)$.
@@
Also, it follows from Corollary 3.17(iii) and (iv) that a $\mathcal{NOS}_{q}(n)$ generated using the approach described in Section 3.3 has asymptotically optimal period as $q\rightarrow\infty$.
@@ 3.6 Construction III: A construction avoiding zeros
It remains to establish how to construct sequences $S$ with the properties of Theorem 3.20. Suppose, for $q>2$ and $n\geq 2$, $G$ is a directed graph with vertices the zero-free $q$-ary $(n-1)$-tuples with weight less than $nq/2-1$, and with a directed edge connecting vertices $\mathbf{u}=(u_{0},u_{1},\ldots,u_{n-2})$ and $\mathbf{v}=(v_{0},v_{1},\ldots,v_{n-2})$ if and only if $u_{i}=v_{i-1}$, $(1\leq i\leq n-2)$, and $w(\mathbf{u})+v_{n-2}<nq/2$. Clearly $G$ is a subgraph of the de Bruijn digraph.
As previously, we identify the edge connecting $(u_{0},u_{1},\ldots,u_{n-2})$ to $(u_{1},u_{2},\ldots,u_{n-2},v)$ with the $n$-tuple $(u_{0},u_{1},\ldots,u_{n-2},v)$. It follows immediately that every zero-free $q$-ary $n$-tuple with weight at most $nq/2$ will appear as an edge in $G$.
@@
must satisfy $t+w(\mathbf{u})<nq/2$. That is, the in-degree of every vertex is the same as its out-degree. Moreover, $G$ is clearly connected, as it is straightforward to construct a ‘low weight’ directed path between any pair of vertices (just as in the discussion in Section 3.3). This means there exists a directed Eulerian Circuit in $G$ (see, for example, [10, Corollary 6.1]), yielding a $q$-ary zero-free $n$-window sequence $S$ with the property that all the $n$-tuples appearing in $S$ have weight less than $nq/2$, and with period equal to the number of zero-free $q$-ary $n$-tuples with weight less than $nq/2$.
@@
There exists a $\mathcal{NOS}_{q}(n)$ of period $(q-1)^{n}/2$ (if $q$ and $n$ are both odd) or $((q-1)^{n}-k_{q,n,nq/2})/2$ (if $q$ or $n$ is even) for all $q>2$ and $n\geq 2$.
Observe that from Theorem 3.10, a $\mathcal{NOS}_{q}(n)$ has period less than $q^{n}/2$. Hence, if $n$ is a fixed odd value, a sequence of $\mathcal{NOS}_{q}(n)$ exist with asymptotically maximal period as $q\rightarrow\infty$ (since, for fixed $n$, $(q-1)^{n}/q^{n}\rightarrow 1$ as $q\rightarrow\infty$).
@@ 3.8.1 Examples for $q=3$
For $q=3$ and small $n$, the zero-free low-weight tuples used in the construction of Section 3.6 are shown in Tables 5 and 6.
@@
These gives rise to a $\mathcal{NOS}_{3}(2)$ of period 1 with ring sequence [1], and a $\mathcal{NOS}_{3}(3)$ of period 4 with ring sequence [1112].
@@ 3.8.2 Examples for $q=4$
For $q=4$ and small $n$, the zero-free low-weight tuples are shown in Tables 7 and 8.
@@
These gives rise to a $\mathcal{NOS}_{4}(2)$ of period 3 with ring sequence [112], and a $\mathcal{NOS}_{4}(3)$ of period 10 with ring sequence [1113112212].
4 Orientable sequences with near-optimal period
We now describe a way of using the negative orientable sequences generated by Construction II to yield orientable sequences of large period for general $n$ and $q$.
4.1 Preliminaries

### v5

@@ 3.3 Construction II: A construction for general $n$
The second method of construction involves the notion of the pseudoweight of an $n$-tuple, which is equal to the sum of the values in a tuple after changing all the zeros to $q/2$. The construction then builds on the observation that if the pseudoweights of all the tuples in a sequence are less than half the maximum possible value ($nq/2$) then the pseudoweight of all the $n$-tuples in the negative sequence will be greater than half the maximum possible value.
We first need the following formal definition of pseudoweight.
@@
As a simple example for $q=3$, the 4-tuple $(0,1,1,2)$ has weight $0+1+1+2=4$ and pseudoweight $1.5+1+1+2=5.5$, since $f(0)=\frac{3}{2}$.
@@
It remains to establish how to construct the desired sequences $S$. Suppose, for $q>2$ and $n\geq 2$, $G$ is a directed graph with vertices the $q$-ary ${(n-1)}$-tuples with pseudoweight less than $nq/2-1$, and with a directed edge connecting vertices $\mathbf{u}=(u_{0},u_{1},\ldots,u_{n-2})$ and $\mathbf{v}=(v_{0},v_{1},\ldots,v_{n-2})$ if and only if $u_{i}=v_{i-1}$, $(1\leq i\leq n-2)$, and $w^{*}(\mathbf{u})+f(v_{n-2})<nq/2$, where $f$ is as in Definition 3.13. As is conventional, we identify the edge connecting $(u_{0},u_{1},\ldots,u_{n-2})$ to $(u_{1},u_{2},\ldots,u_{n-2},v)$ with the $n$-tuple $(u_{0},u_{1},\ldots,u_{n-2},v)$, and $G$ contains as an edge every $q$-ary $n$-tuple with pseudoweight less than $nq/2$.
@@
must satisfy $t+w^{*}(\mathbf{u})<nq/2$. That is, the in-degree of every vertex is the same as its out-degree. Moreover, $G$ is clearly connected, as it is straightforward to construct a ‘low pseudoweight’ directed path between any pair of vertices — that is, for any vertex $\mathbf{u}=(u_{0},u_{1},\ldots,u_{n-2})$ there is a path to $(1,1,\ldots,1)$ via vertices of the form $(u_{i},u_{i+1},\ldots,u_{n-2},1,1,\ldots,1)$ and similarly from $(1,1,\ldots,1)$ to any vertex $\mathbf{v}$. This means there exists a directed Eulerian Circuit in $G$ (see, for example, [11, Corollary 6.1]), yielding a $q$-ary $n$-window sequence $S$ with the property that all the $n$-tuples appearing in $S$ have weight less than $nq/2$, and with period equal to the number of $q$-ary $n$-tuples with pseudoweight less than $nq/2$.
Let $r_{q,n,s}$ ($q\geq 2$, $n\geq 1$) denote the number of $q$-ary $n$-tuples with pseudoweight exactly $s$, where $r_{q,n,s}=0$ by definition if $s<n$ or $s>n(q-1)$. Then the above discussion establishes the following result.
@@
There exists a $\mathcal{NOS}_{q}(n)$ of period $\frac{q^{n}-r_{q,n,nq/2}}{2}$ for all $q>2$ and $n\geq 2$.
@@ 3.4 Enumerating $n$-tuples with given pseudoweight
Lemma 3.15 means that it is of interest to know the value of $r_{q,n,s}$. If $q$ is even then $r_{q,n,s}$ is only defined for integer values $s$, and if $q$ is odd it is only defined when $s$ is a multiple of 0.5. We have the following elementary result.
@@
Suppose $q\geq 2$ and $n\geq 1$.
@@
    Parts (i)–(iv) are trivially true (observing that the 1-tuples $(0)$ and $(q/2)$ both have pseudoweight $q/2$). Part (v) follows immediately by considering the $q$ possible ways of adding a final entry to an $(n-1)$-tuple to make an $n$-tuple of a given pseudoweight, observing that inserting 00 as the final entry will increase the pseudoweight by $q/2$.
We can also make some simple observations.
@@
    If $n\geq 1$ and $0\leq i\leq 2n$ then $r_{3,n,n+i/2}=N_{2}(i,n)$, where $N_{2}(i,n)$ is a trinomial coefficient (see Section 3.7 below).
@@
      If $w\neq q$, then, by the same result, $r_{q,2,w}$ will be the sum of $q$ terms at most $q-1$ of which are non-zero, all equal to 1, i.e. $r_{q,2,w}\leq q-1$.
@@
      If $w\neq q$, then, by the same result, $r_{q,2,w}$ is equal to the sum of $q$ terms at most $q-1$ of which are non-zero, with at most one equal to $2$ and all others no greater than $2$ ($r_{q,n-1,w-q/2}$ is at most one since $w\neq q$). Hence $r_{q,2,w}\leq(q-2)+2=q$.
@@
      This follows by induction on $n$. If $n=3$ then, by Lemma 3.16(v), $r_{q,3,w}$ is the sum of $q$ terms, all of which are less than $q$ except possibly at most one equal to $q$ — from (i). Hence, since $q>2$, $r_{q,3,w}\leq(q-1)^{2}+q<q^{2}$, and the claim holds. The induction step is trivial since $r_{q,m+1,w}$ is equal to the sum of $q$ terms $r_{q,m,i}$, each of which is less than $q^{m-1}$.
@@
      This also follows by induction on $n$. If $n=3$ then, by Lemma 3.16(v), $r_{q,3,w}$ is the sum of $q$ terms, all of which are at most $q$ except possibly at most one equal to $q+2$) — from (ii). Hence, $r_{q,3,w}\leq(q-1)q+(q+2)=q^{2}+2$, and the claim holds. The induction step is trivial since $r_{q,m+1,w}$ is equal to the sum of $q$ terms $r_{q,m,i}$, each of which is at most $q^{m-3}(q^{2}+2)$.
@@
Also, it follows from Corollary 3.17(iii) and (iv) that a $\mathcal{NOS}_{q}(n)$ generated using the approach described in Section 3.3 has, by Lemma 3.15, period at least $(q^{n}-q^{n-3}(q^{2}+2))/2=\frac{q^{3}-q^{2}-2}{2q^{3}}q^{n}$. Since $\frac{q^{3}-q^{2}-2}{q^{3}}$ approaches $1$ as $q\rightarrow\infty$, these negative orientable sequences have period the same order of magnitude as the upper bound. Of course, if $n>2$ the period is unlikely to actually be optimal, as the construction avoids all $n$-tuples of pseudoweight exactly $nq/2$.
The values of $r_{q,n,nq/2}$ for $q=3$, $q=4$ and $1\leq n\leq 6$ are given in Table 2. These values were derived using Corollary 3.17(v) and (vi).
Table 2: Values of $r_{q,n,nq/2}$ for small $q$ and $n$
$n=2$ $n=3$ $n=4$ $n=5$ $n=6$ $n=7$ $n=8$
$q=3$ 3 7 19 51 141 393 1107
$q=4$ 6 20 70 252 924 3432 12870
$q=5$ 5 13 69 261 1301 5685 27525
$q=6$ 8 38 196 1052 5774 32146 180772
$q=7$ 7 19 183 791 6613 35561 267639
$q=8$ 10 62 426 3052 22354 166014 1245066
The periods of the negative orientable sequences using Construction II for $q=3$, $q=4$ and $1\leq n\leq 6$ are given in Table 3. These values were derived using Lemma 3.15 and Table 2.
Table 3: $\mathcal{NOS}_{q}(n)$ periods (Construction II) for small $q$ and $n$
$n=2$ $n=3$ $n=4$ $n=5$ $n=6$ $n=7$ $n=8$
$q=3$ 3 10 31 96 294 897 2727
$q=4$ 5 22 93 386 1586 6476 26333
$q=5$ 10 56 278 1432 7162 36220 181550
$q=6$ 14 89 550 3362 20441 123895 749422
$q=7$ 21 162 1109 8008 55518 393991 2748581
$q=8$ 27 225 1835 14858 119895 965569 7766075
These values can be compared with the bounds given in Table 1.
@@ 3.6 Construction III: A construction avoiding zeros
It remains to establish how to construct sequences $S$ with the properties of Theorem 3.20. Suppose, for $q>2$ and $n\geq 2$, $G$ is a directed graph with vertices the zero-free $q$-ary $(n-1)$-tuples with weight less than $nq/2-1$, and with a directed edge connecting vertices $\mathbf{u}=(u_{0},u_{1},\ldots,u_{n-2})$ and $\mathbf{v}=(v_{0},v_{1},\ldots,v_{n-2})$ if and only if $u_{i}=v_{i-1}$, $(1\leq i\leq n-2)$, and $w(\mathbf{u})+v_{n-2}<nq/2$.
As previously, we identify the edge connecting $(u_{0},u_{1},\ldots,u_{n-2})$ to $(u_{1},u_{2},\ldots,u_{n-2},v)$ with the $n$-tuple $(u_{0},u_{1},\ldots,u_{n-2},v)$. It follows immediately that every zero-free $q$-ary $n$-tuple with weight less than $nq/2$ will appear as an edge in $G$.
@@
must satisfy $t+w(\mathbf{u})<nq/2$. That is, the in-degree of every vertex is the same as its out-degree. Moreover, $G$ is clearly connected, as it is straightforward to construct a ‘low weight’ directed path between any pair of vertices (just as in the discussion in Section 3.3). This means there exists a directed Eulerian Circuit in $G$ (see, for example, [11, Corollary 6.1]), yielding a $q$-ary zero-free $n$-window sequence $S$ with the property that all the $n$-tuples appearing in $S$ have weight less than $nq/2$, and with period equal to the number of zero-free $q$-ary $n$-tuples with weight less than $nq/2$.
@@
There exists a $\mathcal{NOS}_{q}(n)$ containing no zeros of period $(q-1)^{n}/2$ (if $q$ and $n$ are both odd) or $((q-1)^{n}-k_{q,n,nq/2})/2$ (if $q$ or $n$ is even) for all $q>2$ and $n\geq 2$.
We show, in the next subsection, that for $q>2$, $k_{q,n,nq/2}<(q-1)^{n-1}$. Thus a sequence of $\mathcal{NOS}_{q}(n)$ exists with period at least $\frac{q-2}{q-1}(q-1)^{n}/2$. Although this period is not of the same order of magnitude as the upper bound of Theorem 3.10 we note that the sequences constructed in this section are negative orientable sequences of order $n$ over the alphabet $\mathbb{Z}_{q}\backslash\{0\}$ of size $q-1$ and the period of such sequences is bounded above by $(q-1)^{n}/2$. Now $\frac{q-2}{q-1}\geq 1/2$ and approaches $1$ as $q\rightarrow\infty$ so the period is the same order of magnitude as this upper bound. Indeed when $nq$ is odd this upper bound is attained.
@@
The values of $k_{q,n,nq/2}$ for $q=3$, $q=4$ and $1\leq n\leq 6$ are given in Table 7 (observing that $k_{q,n,nq/2}$ is undefined for $q$ and $n$ both odd). These values were derived using Lemma 3.23.
Table 7: Values of $k_{q,n,nq/2}$ for $q=3$ and $q=4$
$n=2$ $n=3$ $n=4$ $n=5$ $n=6$
$q=3$ 2 6 20
$q=4$ 3 7 19 51 141
The periods of the negative orientable sequences using Construction III for $q=3$, $q=4$ and $1\leq n\leq 6$ are given in Table 8. These values were derived using Lemma 3.21 and Table 7.
Table 8: $\mathcal{NOS}_{q}(n)$ periods (Construction III) for $q=3$ and $q=4$
$n=2$ $n=3$ $n=4$ $n=5$ $n=6$
$q=3$ 1 4 5 16 22
$q=4$ 3 10 31 96 294
@@ 3.7 Enumerating $n$-tuples of given weight
Corollary 3.24
Suppose $q>2$.
  1. (i)
    If $22\leq w\leq 2q-$ 2, then $k_{q,2,q}=q-1$ and $k_{q,2,w}<q-1$ if $w\neq q$.
  2. (ii)
    If $n\geq 3$ and $n\leq w\leq n(q-1)$, then $k_{q,n,w}<(q-1)^{n-1}$.
  • Proof
    • (i)
      This follows immediately by Lemma 3.23(iv)).
    • (ii)
      By Lemma 3.23(iii),
[…]
      The result now follows by induction using Lemma 3.23(iii).
@@ 3.8.1 Examples for $q=3$
For $q=3$ and small $n$, the zero-free low-weight tuples used in the construction of Section 3.6 are shown in Tables 9 and 10.
@@
The single 2-tuple in Table 9 and the four 3-tuples in Table 10 give rise to a $\mathcal{NOS}_{3}(2)$ of period 1 with ring sequence [1], and a $\mathcal{NOS}_{3}(3)$ of period 4 with ring sequence [1112], respectively.
@@ 3.8.2 Examples for $q=4$
For $q=4$ and small $n$, the zero-free low-weight tuples are shown in Tables 11 and 12.
@@
The three 2-tuples in Table 11 and the ten 3-tuples in Table 12 give rise to a $\mathcal{NOS}_{4}(2)$ of period 3 with ring sequence [112], and a $\mathcal{NOS}_{4}(3)$ of period 10 with ring sequence [1113112212], respectively.
4 Orientable sequences
We now describe a way of using the negative orientable sequences generated by Construction II to yield orientable sequences for general $n$ and $q$ with period of the same order of magnitude as the upper bound. We first need to introduce the Lempel Homorphism, which is fundamental to the construction of orientable sequences.
4.1 The de Bruijn graph and the Lempel Homomorphism
Definition 4.1
For a nonzero $\beta\in\mathbb{Z}_{q}$, we define a function $D_{\beta}$ from $B_{n}(q)$ to $B_{n-1}(q)$ as follows. For $\textbf{a}=(a_{1},\dots,a_{n})$ and $\textbf{b}=(b_{1},\dots,b_{n-1}),\ D_{\beta}(\textbf{a})=\textbf{b}$ if and only if $b_{i}=d_{\beta}(a_{i},a_{i+1})$ for $i=1$ to $n-1$, where $d_{\beta}(a_{i},a_{i+1})=\beta(a_{i+1}-a_{i})\mod q.$
$D((s_{i}))=\{(t_{i}):t_{j}=\beta(s_{j+1}-s_{j})\}.$
We are particularly interested in the case $\beta=1$, and we simply write $D$ for $D_{1}$. Examples of the application of $D^{-1}$ are given in Sections 4.5 and 5.3.
The weight $w(S)$ of a cycle $S$ is the weight of the ring sequence corresponding to $S$ (that is the sum of the terms $s_{0},s_{1},\dots,s_{m-1}$ treating $s_{i}$ as an integer in the range $[0,q-1]$). Similarly we write $w_{q}(S)$ for $w(S)\bmod{q}$.
4.2 Construction methods